# EE-MLLM: A Data-Efficient and Compute-Efficient Multimodal Large Language Model

### v1

@@ Introduction
where $X^{q}=[I;T]W_{q}$, $X^{k}=[I;T]W_{k}$, $X^{v}=[I;T]W_{v}$, $W_{q}$, $W_{k}$ and $W_{v}$ are the weights of the self-attention module. The symbol [;][;][ ; ] denotes the concatenation operation of two variables in the sequence dimension. As illustrated in Fig. 2, our composite attention module uses text tokens as the queries and the concatenation of visual tokens and text tokens as keys and values for cross-attention. Specifically, $X^{q}=TW_{q}$, $X^{k}=[I;T]W_{k}$, and $X^{v}=[I;T]W_{v}$. The length of the output sequence will be $n$, which is the same as the length of text tokens $T$. The attention mask for self-attention is a lower triangular matrix, whereas the attention mask for the composite attention module is a trapezoidal matrix. This implies that text tokens will attend to all previous tokens, including visual tokens, while the interaction within visual tokens is eliminated.
@@
Aligner. We introduce the aligner, which leverages the existing weights at each layer of LLM to enhance modality alignment between vision and language without the need for additional modules such as cross-attention-based methods. By reusing these weights, visual tokens aligned in the input space of the LLM can naturally facilitate modality alignment at each layer. This approach eliminates the necessity for the training procedure to concentrate on modality alignment within the LLM, thereby promoting data efficiency. As depicted in Fig. 2, in each layer, text tokens are mapped by the value matrix ($W_{v}$), output matrix ($W_{o}$), and FFN. Hence, we make visual tokens follow the same mapping relationship in the aligner. Specifically, the aligner comprises $W_{v}$, $W_{o}$, and FFN, which share weights with the corresponding LLM modules. The residual connection is also implemented in the aligner. The aligner can be formulated as follows:
$H$ $=IW_{v}W_{o}+I$ (2)
$O$ $={\rm FFN}(H)+H.$ (3)
This can be viewed as a decoder block without attention.
Analysis of the Computational Overhead
Our EE-MLLM can reduce the computational overhead without introducing additional parameters. In this section, we analyze the compute efficiency of EE-MLLM from the perspective of FLOPs, comparing it with LLaVA. The LLM consists of $d$ blocks, with a hidden states dimension of $h$, an input text length of $T$, and $V$ visual tokens. Considering an LLM, we assume the input sequence length is $L$, the FLOPs of the self-attention module and FFN are $8Lh^{2}+4L^{2}h$ and $16Lh^{2}$. For self-attention-based methods, such as LLaVA, the length of the input sequence is $L=V+T$. Therefore, the FLOPs of LLaVA are $24(T+V)dh^{2}+4(T+V)^{2}dh$. For EE-MLLM, the FLOPs of composite attention modules are $(6T+2V)h^{2}+4VTh+4T^{2}h$. Given that we introduce aligners, the FLOPs of FFN are the same as those of LLaVA, which are $16(V+T)h^{2}$. The aligners introduce an additional matrix multiplication by $O$, adding $2Vh^{2}$ FLOPs. Therefore, the total FLOPs of EE-MLLM are $2(11T+10V)dh^{2}+4VTdh+4T^{2}dh$. Particularly, when the resolution of the input image is $980\times 980$, the number of visual tokens is 4,900, and the length of text tokens is 256, which is a typical length during training, the FLOPs of EE-MLLM amount to 70% of those of LLaVA.
Experiments
Implementation Details
Model Configurations. We employ Vicuna-7b-v1.5 (Zheng et al. 2023) as our LLM and SigLIP (Zhai et al. 2023) as the vision encoder. Specifically, SigLIP is initialized from Idefics2 (Laurençon et al. 2024), which supports dynamic resolutions up to 980 $\times$ 980. The projector consists of a two-layer MLP, identical to that of LLaVA (Liu et al. 2023b).
Model LLM MMB-T MME ScienceQA HallB MMMU CCBench SeedBench BLINK
InstructBLIP Vicuna-7B 36.0 1137.1 54.7 31.2 30.6 12.7 44.5 -
MiniGPT-4-v1 Vicuna-7B 12.2 770.6 39.0 31.9 23.6 1.8 31.6 -
MiniGPT-4-v2 Llama2-7B 24.3 708.4 54.1 30.0 25.0 1.4 29.4 -
Idefics-instruct Llama-7B 48.2 942 51.6 27.3 18.4 7.8 45.0 38.3
OpenFlamingo v2 MPT-7B 6.6 535 45.7 29.4 28.2 6.3 28.8 -
Qwen-VL Qwen-7B 38.2 334.1 57.7 29.9 29.6 6.1 52.5 27.9
Qwen-VL-Chat Qwen-7B 60.6 1467.8 65.5 36.8 37.0 41.2 64.8 28.2
VLoRA Vicuna-7B 63.4 1311.3 66.4 26.4 36.0 28.6 - -
ShareGPT4V Vicuna-7B 64.6 1561.4 68.2 28.6 37.2 30.8 69.3 40.9
LLaVA-v1.5 Vicuna-7B 62.3 1510.7 66.8 27.6 35.7 27.5 65.8 39.7
LLaVA-v1.6 Vicuna-7B 66.5 1475.6 68.5 27.6 37.6 24.3 69.6 41.6
EE-MLLM Vicuna-7B 70.4 1528.1 77.7 38.6 33.4 37.3 70.2 43.2
Table 2: Comparsion with state-of-the-art methods on general benchmarks, including MMBench, MME, ScienceQA, HallusionBench, MMMU, CCBench, SeedBench, and BLINK. MMB-T indicates we report the results on MMBench-TEST-EN-V11.
Pre-training Configurations. During the pre-training stage, we follow LLaVA to adopt blip-558k (Li et al. 2022) for training one epoch. We freeze the LLM and the vision encoder, making only the projector trainable. We adopt AdamW (Loshchilov and Hutter 2018) optimizer. The learning rate is 1e-3, followed by a linear warm-up scheduler and then a cosine decay scheduler. The global batch size is 256.
Fine-tuning Configurations. During the fine-tuning stage, we freeze the vision encoder and update the weights of LLM and projector. The total quantity of supervised fine-tuning data amounts to 3 million. As displayed in Tab. 1, following Deepseek-VL (Lu et al. 2024), our supervised fine-tuning data incorporates open-source gpt4v datasets, including ShareGPT4V (Chen et al. 2023b), LVIS-Instruct4V (Wang et al. 2023a), LAION-GPT4V (LAION 2023), and TextOCR-GPT4V (Carter 2024). Moreover, to enhance the fine-grained capabilities of MLLM, we introduce DVQA (Kafle et al. 2018), PlotQA (Methani et al. 2020), CLEVR (Johnson et al. 2017), DocVQA (Mathew, Karatzas, and Jawahar 2021), ChartQA (Masry et al. 2022), and ScienceQA (Lu et al. 2022) for fine-tuning. To preserve language abilities during fine-tuning, we utilize a large volume of text-only data from various sources. The text-only data encompasses various dataset types, such as mathematical reasoning, multi-round dialogues, logic comprehension, and code. Meanwhile, the in-house data primarily consists of OCR-related, table and diagram comprehension, and mathematical reasoning data. The optimizer is AdamW with a global batch size of 128. The learning rate is 2e-5 and the learning rate scheduler is the same as the pre-training stage. The training is conducted on 16 NVIDIA H800 GPUs.
Evalution Benchmarks

### v3

@@ 3.1.1 Composite Attention Module
As illustrated in Fig. 2, our composite attention module uses text tokens as queries and the concatenation of visual tokens and text tokens as keys and values for cross-attention. Specifically, $X^{q}=TW_{q}$, $X^{k}=[I;T]W_{k}$, and $X^{v}=[I;T]W_{v}$. The length of the output sequence will be $n$, which is the same as the length of text tokens $T$. The attention mask for self-attention is a lower triangular matrix, whereas the attention mask for the composite attention module is a trapezoidal matrix. This implies that text tokens will attend to all previous tokens, including visual tokens, while intra-interaction among visual tokens is eliminated.
3.1.2 Aligner
We introduce the aligner, which leverages the existing weights at each layer of LLM to enhance modality alignment between vision and language without the need for additional modules such as cross-attention-based methods. By reusing these weights, visual tokens aligned in the input space of the LLM can naturally facilitate modality alignment at each layer. This approach eliminates the necessity for the training procedure to concentrate on modality alignment within the LLM, thereby promoting data efficiency.
As depicted in Fig. 2, in each layer, text tokens are mapped by the value matrix ($W_{v}$), output matrix ($W_{o}$), and FFN. Hence, we make visual tokens follow the same mapping relationship in the aligner. Specifically, the aligner comprises $W_{v}$, $W_{o}$, and FFN, which share weights with the corresponding LLM modules. The residual connection is also implemented in the aligner. The aligner can be formulated as follows:
$O=\mathrm{FFN}(IW_{v}W_{o}+I)+IW_{v}W_{o}+I$ (2)
This part of the decoder layer skips attention calculation.
3.2 Training-Free Adaptation of EE-MLLM
Since our composite attention mechanism does not introduce additional parameters for the interaction between vision and language, which reuses the weights of each LLM layer for modality alignment, we can directly apply this method to self-attention-based models without requiring additional training. Specifically, given a well-trained MLLM with $L$ layers, we can replace the self-attention blocks in selected layers $\mathcal{L}_{\text{replace}}\subseteq\{1,2,...,L\}$ with our composite attention. For each layer $l\in\mathcal{L}_{\text{replace}}$, the original self-attention is substituted with our composite attention, and the weights of $W_{v}$, $W_{o}$, and FFN are reused to align the visual tokens.
Inspired by FastV [9], which demonstrates the attention of visual tokens becomes sparse in deep layers, we maintain the self-attention mechanism in the early layers of LLM while replacing the deeper layers with our composite attention. With this adaption, we can reduce the computation cost of the self-attention-based MLLM in deep layers, while maintaining strong performance.
Model LLM MMB-T MME ScienceQA HallB MMMU CCBench SeedBench BLINK
InstructBLIP Vicuna-7B 36.0 1137.1 54.7 31.2 30.6 12.7 44.5
MiniGPT-4-v1 Vicuna-7B 12.2 770.6 39.0 31.9 23.6 1.8 31.6
MiniGPT-4-v2 Llama2-7B 24.3 708.4 54.1 30.0 25.0 1.4 29.4
Idefics-instruct Llama-7B 48.2 942.0 51.6 27.3 18.4 7.8 45.0 38.3
OpenFlamingo v2 MPT-7B 6.6 535.0 45.7 29.4 28.2 6.3 28.8
Qwen-VL Qwen-7B 38.2 334.1 57.7 29.9 29.6 6.1 52.5 27.9
Qwen-VL-Chat Qwen-7B 60.6 1467.8 65.5 36.8 37.0 41.2 64.8 28.2
VLoRA Vicuna-7B 63.4 1311.3 66.4 26.4 36.0 28.6
ShareGPT4V Vicuna-7B 64.6 1561.4 68.2 28.6 37.2 30.8 69.3 40.9
LLaVA-v1.5 Vicuna-7B 62.3 1510.7 66.8 27.6 35.7 27.5 65.8 39.7
LLaVA-v1.6 Vicuna-7B 66.5 1475.6 68.5 27.6 37.6 24.3 69.6 41.6
EE-MLLM Vicuna-7B 70.4 1528.1 77.7 38.6 33.4 37.3 70.2 43.2
Table 1: Comparison with state-of-the-art methods on general benchmarks, including MMBench, MME, ScienceQA, HallusionBench, MMMU, CCBench, SeedBench, and BLINK. MMB-T indicates we report the results on MMBench-TEST-EN-V11.
Model LLM AI2D OCRBench TextVQA ChartQA Seed2 Plus
InstructBLIP Vicuna-7B 40.6 27.6 33.6 10.9 29.5
MiniGPT-4-v1 Vicuna-7B 28.4 17.2 15.2
MiniGPT-4-v2 Llama2-7B 30.5 3.1 23.3
Idefics-instruct Llama-7B 42.2 25.2 35.4
OpenFlamingo v2 MPT-7B 31.7 14.9 16.3 28.7
Qwen-VL Qwen-7B 57.7 12.7 63.1 59.0 40.1
Qwen-VL-Chat Qwen-7B 63.0 48.8 60.7 49.8 40.6
ShareGPT4V Vicuna-7B 58.0 37.1 51.1 21.3 46.1
LLaVA-v1.5 Vicuna-7B 55.5 31.8 45.5 17.8 41.3
LLaVA-v1.6 Vicuna-7B 67.0 53.2 64.4 55.4 51.6
EE-MLLM Vicuna-7B 65.7 57.4 69.0 68.4 53.5
Table 2: Comparison with state-of-the-art methods on fine-grained benchmarks, including AI2D, OCRBench, TextVQA, ChartQA, and SeedBench-2 Plus.
3.3 Analysis of the Computational Overhead
EE-MLLM reduces computational overhead without introducing extra parameters. This section analyzes the compute efficiency of EE-MLLM compared to LLaVA in terms of FLOPs. We consider an LLM with $d$ blocks, hidden state dimension $h$, intermediate size $m$, $T$ text tokens, and $V$ visual tokens.
The total FLOPs of LLaVA are given by:
$\text{FLOPs}_{\text{LLaVA}}=\underbrace{8(V+T)dh^{2}+4(V+T)^{2}dh}_{\text{self-attention module}}+\underbrace{6(V+T)dhm}_{\text{FFN module}}.$ (3)
The total FLOPs of EE-MLLM are given by:
$\begin{split}\text{FLOPs}_{\text{EE-MLLM}}&=\underbrace{(6V+8T)dh^{2}+(2T^{2}+2VT)dh}_{\text{composite attention module}}\\ &\quad+\underbrace{6(V+T)dhm}_{\text{FFN module}}.\end{split}$ (4)
For a typical training scenario with a $980\times 980$ input image (yielding 4,900 visual tokens) and 256 text tokens, EE-MLLM reduces computational requirements by approximately 24% compared to LLaVA, demonstrating significant efficiency gains.
4 Experiments
4.1 Implementation Details
Model Configurations. We employ Vicuna-7b-v1.5 [47] as our LLM and SigLIP [46] as the vision encoder. Specifically, SigLIP is initialized from Idefics2 [21], which supports dynamic resolutions up to 980 $\times$ 980. The projector consists of a two-layer MLP.
Pre-training Configurations. During the pre-training stage, we follow LLaVA to adopt blip-558k [25] for training one epoch. We freeze the LLM and the vision encoder, making only the projector trainable. We adopt AdamW [32] optimizer. The learning rate is 1e-3, followed by a linear warm-up scheduler and then a cosine decay scheduler. The global batch size is 256.
Fine-tuning Configurations. During the fine-tuning stage, we freeze the vision encoder and update the weights of the LLM and projector. The total amount of supervised fine-tuning data is 3 million samples. The sources of fine-tuning data are provided in Appendix 8.1. We utilize the AdamW optimizer with a global batch size of 128 and a learning rate of 2e-5. The learning rate scheduler is the same as the pre-training stage. Training is conducted on 16 H800 GPUs.
4.2 Evalution Benchmarks
We conduct our evaluations using the VLMEvalKit [11], and the results of other state-of-the-art models are obtained from the same source.
General Benchmarks. For general benchmarks, our evaluation includes MMBench-EN [30], MME [13], ScienceQA [34], HallusionBench [15], MMMU [45], CCBench [30], SeedBench [23], and BLINK [14]. Among them, MMBench-EN is a comprehensive multimodal benchmark specifically designed to assess the performance of multimodal large language models (MLLMs). It contains more than 3,000 multiple-choice questions spanning 20 distinct ability categories. We evaluate EE-MLLM on MMBench-EN-V1.1. Further details on the other general benchmarks are provided in Appendix 9.1.
Fine-grained Benchmarks. We also compare EE-MLLM with other models on AI2D [19], OCRBench [31], TextVQA [40], ChartQA [36], DocVQA [37], and SeedBench 2 Plus [24]. More details of these fine-grained benchmarks are provided in Appendix 9.2.
Figure 3: Comparison of prefilling time between EE-MLLM and LLaVA. The X-axis represents the number of input images, where each image has a resolution of $980\times 980$. The Y-axis indicates the prefilling time.
4.3 Comparisons with State-of-the-art Models
General Benchmarks. In Tab. 1, we compare EE-MLLM with various MLLMs equipped similar LLM [29, 3, 35] on general benchmarks. These benchmarks evaluate the comprehensive ability of MLLMs, including understanding and perception. EE-MLLM achieves comparable performance with state-of-the-art MLLMs in general benchmarks. Specifically, EE-MLLM achieves a score of 70.4 in MMBench and 1528.1 in MME, scores that are notably higer compared to LLaVA-v1.6, which also support high-resolution images input. This suggests that EE-MLLM possesses comprehensive perception and reasoning capabilities. What’s more, EE-MLLM also achieves promising performance on CCBench and SeedBench.
Fine-grained Benchmarks. In Tab. 2, we evaluate EE-MLLM on fine-grained benchmarks. These benchmarks demand strong visual perception capabilities from MLLMs, as they require extracting fine-grained information from images to accurately answer questions. On TextVQA and ChartQA, EE-MLLM achieves impressive performance, 4.6 higher on TextVQA and 13.0 higher on ChartQA compared with LLaVA-v1.6, which also supports high resolution. In the OCRBench, specifically designed to assess the OCR capabilities of MLLM, EE-MLLM outperforms by 4.2 compared to LLaVA-v1.6. These results demonstrate that despite EE-MLLM significantly reducing the computational overhead associated with visual tokens, it effectively maintains the model’s fine-grained capabilities.
Model MMB-VAL SeedBench ScienceQA HallB
LLaVA-v1.5 67.3 65.1 67.0 25.7
EE-MLLM-F 67.4 65.1 67.0 32.4
AI2D OCRBench TextVQA ChartQA
LLaVA-v1.5 56.4 31.6 46.1 17.9
EE-MLLM-F 56.4 30.7 45.5 17.4
Table 3: Performance comparison of LLaVA-v1.5 with our composite attention mechanism across general and fine-grained benchmarks.
4.4 Comparison of Prefilling Time
Since we eliminate the intra-interactions among visual tokens, EE-MLLM can significantly reduce the prefilling time, thus accelerating the generation of the first output token. This efficiency is crucial in real-time scenarios, particularly for tasks that only require a few output tokens. For example, in content moderation systems, MLLMs must rapidly classify social medias as safe or unsafe quickly to prevent harmful content exposure while maintaining platform throughput. Similarly, industrial quality control systems require MLLMs to swiftly provide pass/fail decisions. We conduct comparisons of prefilling time on a single NVIDIA H800. The resolution of the input image is set to $980\times 980$. We illustrate the time consumed for prefilling by EE-MLLM and LLaVA in Fig. 3.
With one input image, EE-MLLM costs 79 ms for prefilling, while LLaVA costs 277 ms. When the input is extended to two images, the time consumed by LLaVA severely increases to 1086 ms, but the prefilling time of EE-MLLM only rises to 158 ms. We also test the scenario where the number of input images is 3, and the gap between EE-MLLM and LLaVA becomes even larger.
To further analyze the efficiency of EE-MLLM, we compare the overall inference speed and provide a detailed analysis in Appendix 10.
4.5 Training-free Adaption of EE-MLLM
In Tab 3, we present the results of directly applying our composite attention mechanism to LLaVA-v1.5-7b without additional training. Specifically, we replace the attention mechanism in the last 16 layers of LLaVA-v1.5-7b with our composite attention approach. And EE-MLLM-F performs comparably to the original LLaVA-v1.5 on MMBench, SeedBench, and ScienceQA, and even achieves better results on HallusionBench. For fine-grained benchmarks, we observe only a slight performance drop on OCRBench and TextVQA. These results demonstrate that our composite attention mechanism can directly adapt to self-attention-based methods without the need for training, while maintaining performance without significant degradation.
Model MMB-VAL SeedBench ScienceQA HallB AI2D OCRBench TextVQA ChartQA Avg. Ratio
Resolution: $336\times 336$
LLaVA-v1.5 67.3 65.1 67.0 25.7 56.4 31.6 46.1 17.9 47.1 100%
Flamingo 60.7 53.0 66.9 28.2 52.2 13.0 27.4 12.6 39.3 81.7%
EE-MLLM 66.8 64.1 66.7 25.2 56.2 29.1 43.3 17.3 46.1 97.9%
Resolution: $672\times 672$
LLaVA-v1.5 70.1 68.0 69.4 27.2 58.3 39.8 61.7 27.9 52.8 100%
Flamingo 64.2 55.6 69.4 28.1 53.9 16.5 36.7 19.8 43.0 81.4%
EE-MLLM 68.2 67.3 69.1 25.9 57.5 37.0 60.1 25.1 51.3 97.2%
Table 4: Comparisons with LLaVA and Flamingo under the same settings, including LLM, vision encoder and training data.
5 Ablation Study
5.1 Implementation Details
Following LLaVA-v1.5 [27], we employ Vicuna-7b-v1.5 [47] as our LLM, and CLIP-ViT-L-14 as vision encoder. The training data is consistent with LLaVA-v1.5. To support high-resolution image input, we replace CLIP-ViT-L-14 with SigLIP [46], which is derived from Idefics2 [21] and supports high-resolution image input. To compare with cross-attention based methods, we reproduce Flamingo under the similar settings, the implementation details are provided in Appendix 8.2.
5.2 Comparisons with LLaVA and Flamingo
In Tab. 4, we reproduce LLaVA and Flamingo under the same settings with EE-MLLM. We follow LLaVA to employ blip-558k for pre-training and adopt LLaVA-665k for fine-tuning. For the resolution of $336\times 336$, we employ CLIP-ViT-L-14 as the vision encoder, ensuring complete alignment with LLaVA-v1.5. EE-MLLM demonstrates comparable performance to LLaVA on general benchmarks. In terms of fine-grained benchmarks, at $336\times 336$ resolution, EE-MLLM exhibits promising performance on AI2D and ChartQA but slightly underperforms LLaVA on OCRBench and TextVQA. The average score for the $336\times 336$ resolution is 46.1, which is 97.9% of LLaVA’s 47.1. And the performance of Flamingo is significantly worse than that of EE-MLLM and LLaVA-v1.5, with an average score of only 39.3, which is 81.7% of LLaVA’s score, this underperformance can be attributed to Flamingo’s architecture. Flamingo introduces additional learnable parameters and requires alignment of visual tokens across different LLM layers, necessitating a large volume of training data, for which the BLIP-558K and LLaVA-665K datasets are insufficient. For the resolution of $672\times 672$, we utilize SigLIP as the vision encoder. As shown in Tab. 4, EE-MLLM obtains comparable results on AI2D and TextVQA, and the average score maintains 97.2% of LLaVA’s performance. Similarly, Flamingo only achieves an average score of 43.0, which can be attributed to inadequate training data.
Aligner Components Evaluation Benchmarks
V O FFN SeedB Hall TextVQA ChartQA DocVQA Avg.
Resolution: $336\times 336$
65.3 24.4 44.8 14.4 18.4 33.5
66.3 25.9 45.5 16.7 19.2 34.7
65.4 25.3 45.4 15.6 19.1 34.2
65.8 25.1 43.5 14.8 17.9 33.4
66.1 25.7 46.1 15.8 18.6 34.5
Resolution: $672\times 672$
66.1 25.2 56.8 19.7 28.6 39.3
66.6 25.5 58.2 23.0 29.8 40.6
66.1 24.3 58.2 21.6 30.1 40.1
66.4 24.1 56.3 18.5 26.5 38.4
66.9 27.6 58.8 24.0 31.8 41.8
Table 5: Ablation study of the aligner components V, O, and FFN. Checkmarks indicate employed weights. To reduce ablation costs, we adopt SigLIP with dynamic resolution from Idefics2, enabling direct evaluation at $672\times 672$ from models trained at $336\times 336$.
Figure 4: The visualization results of examples sampled from BLINK and RealWorldQA. Answers are generated by our EE-MLLM.
5.3 Ablation Study of the Aligner
In Tab. 5, we remove the different weights from the aligner and train the model at a resolution of $336\times 336$. Notably, to reduce the cost of ablation experiments, we adopt SigLIP with dynamic resolution support, derived from Idefics2, as our vision encoder. Thus, models trained at $336\times 336$ can be directly evaluated at a resolution of $672\times 672$.
We have three findings:
1) Removing the entire aligner, as depicted in the first row, significantly reduces performance across multiple benchmarks. Specifically, the TextVQA score drops from 46.1 to 44.8, and the average score decreases from 34.5 to 33.5. This result highlights the aligner’s effectiveness in aligning visual features with the LLM feature space, allowing text tokens to capture crucial visual information and address questions through causal cross-attention modules.
2) When ablating individual weights in the aligner, we find that maintaining the structure is more critical. The absence of $V$ or $O$ has a relatively minor impact on low-resolution input, even resulting in slightly higher average performance when $V$ is missing. However, when the FFN is missing, the aligner’s structure no longer resembles a transformer block, leading to a significant performance loss.
3) The importance of a complete aligner becomes even more pronounced at higher resolution ($672\times 672$). When $V$ or $O$ weights are removed, we observe substantial performance drops on detail-oriented benchmarks like TextVQA, with scores decreasing from 58.8 to 58.2, ChartQA from 24.0 to 21.6, and DocVQA reducing from 31.8 to 30.1. This demonstrates that all components are essential for effectively processing high-resolution visual inputs, where fine-grained details significantly impact performance.
5.4 Ablation Study of Training-free Adaption
In Tab. 6, we explore different layer configurations for training-free adaptation of EE-MLLM. Using LLaVA-v1.5-7b with 32 layers as our baseline, we investigate various configurations of attention mechanism modification.
The first configuration represents the original LLaVA, where all layers from 0 to 31 use the standard self-attention mechanism, achieving an average score of 35.5. When we apply the composite attention mechanism to layers 8 to 31, we observe a slight performance decrease to 32.4 on average, with improvements only on HallusionBench.
Notably, the third configuration, which employs composite attention for layers 16 through 31 while maintaining self-attention for layers 0 through 15, achieves the best overall performance with a 36.5 average. This configuration not only maintains comparable performance on SeedBench, TextVQA, ChartQA, and DocVQA compared to the baseline, but also demonstrates significant improvement on HallucinationBench with a score of 32.4 versus 25.7.
To further validate the effectiveness of our composite attention mechanism, we implement a configuration in the fourth row that maintains self-attention only in layers 0 through 15 and drops visual tokens in deeper layers. This results in a substantial performance degradation across fine-grained benchmarks including TextVQA, ChartQA, and DocVQA, with the average score dropping to 27.2. This finding confirms the composite attention mechanism plays a crucial role in preserving performance on tasks requiring detailed visual understanding, while simply dropping visual information in deeper layers is detrimental.
These experimental findings clearly suggest that applying composite attention selectively to deeper layers from 16 to 31 provides an optimal balance between efficiency and overall performance across diverse multimodal tasks.
Layer Configuration Evaluation Benchmarks
Self-attn Layers Comp-attn Layers SeedB Hall TextVQA ChartQA DocVQA Avg.
$[0,31]$ 65.1 25.7 46.1 17.9 22.5 35.5
$[0,7]$ $[8,31]$ 62.8 29.0 39.8 16.1 14.5 32.4
$[0,15]$ $[16,31]$ 65.1 32.4 45.5 17.4 21.9 36.5
$[0,15]$ 65.1 29.2 17.6 13.8 10.5 27.2
Table 6: Ablation study on layer configuration for training-free adaptation. “Self-attn Layers” indicates which layers maintain the original self-attention mechanism, while “Comp-attn Layers” shows which layers are modified to use the composite attention mechanism. The notation $[i,j]$ indicates layers from index $i$ to $j$.
5.5 Visualization
We present four examples sampled from BLINK [14] and RealWorldQA [43] to assess the impact of the architectural changes in Fig 4. The first example demonstrates that EE-MLLM can perceive fine-grained visual context within the image, such as the color of traffic lights. The second and third examples highlight EE-MLLM’s ability to comprehend the position of objects. Specifically, EE-MLLM can accurately identify guitar’s position.
6 Conclusion
In this paper, we introduce EE-MLLM, a novel framework that unifies data efficiency and compute efficiency through two key innovations: (1) a composite attention mechanism that eliminates redundant self-attention computations among visual tokens, reducing FLOPs by 24% compared to self-attention methods, and (2) a parameter-free aligner that reuses existing LLM weights to align visual and text modalities, minimizing training data requirements. EE-MLLM achieves state-of-the-art performance across various benchmarks. Crucially, it reduces inference latency from 277 ms to 79 ms, enabling real-time applications such as content moderation and quality control. The training-free variant (EE-MLLM-F) further demonstrates adaptability, maintaining performance while cutting computational costs by selectively applying composite attention to deeper layers.
Supplementary Material
Figure 5: Comparison of EE-MLLM and LLaVA-v1.6 response results on challenging samples.
7 Visualization
Fig. 5 provides two examples comparing EE-MLLM with LLaVA-v1.6 [29]. The first image, a handwritten coffee menu, prompts a question about the image’s fine-grained details. Our EE-MLLM successfully identifies the price of the Flat White listed in the image, providing the correct answer, whereas LLaVA-v1.6 does not. The second example requires the model to convert the image’s table into markdown format. Our observations indicate that EE-MLLM’s response is more accurate, while LLaVA makes mistakes in the columns labeled “Open Source” and “Data Source”.
8 Implementation Details
8.1 Supervised Fine-tuning Data
The details of our supervised fine-tuning data are shown in Tab. 7 Following Deepseek-VL [33], our supervised fine-tuning data incorporates open-source gpt4v datasets, including ShareGPT4V [8], LVIS-Instruct4V [41], LAION-GPT4V [20], and TextOCR-GPT4V [4]. Moreover, to enhance the fine-grained capabilities of MLLM, we introduce DVQA [18], PlotQA [38], CLEVR [17], DocVQA [37], ChartQA [36], and ScienceQA [34] for fine-tuning. To preserve language abilities during fine-tuning, we utilize a large volume of text-only data from various sources. The text-only data encompasses various dataset types, such as mathematical reasoning, multi-round dialogues, logic comprehension, and code. Meanwhile, the in-house data primarily consists of OCR-related, table and diagram comprehension, and mathematical reasoning data.
8.2 Implementation of Flamingo
In the ablation study, we reproduced Flamingo. Like LLaVA-v1.5-7B, we used CLIP-ViT-L-14 as the vision encoder and Vicuna-7B as the LLM. We introduced cross-attention blocks every 4 layers into the LLM for inter-modal interaction. The cross-attention block includes attention operations and MLP. Specifically, we inserted cross-attention blocks at layers 3, 7, 11, 15, 19, 23, 27, and 31. The training setup was consistent with LLaVA, including training data, learning rate, and optimizer. In the ablation, we ensured a fair comparison.
@@ 8.2 Implementation of Flamingo
9 Evalution Benchmarks